### v1

# SLOctolyzer: Fully automatic analysis toolkit for segmentation and feature extracting in scanning laser ophthalmoscopy images

Abstract
Purpose To describe SLOctolyzer: an open-source analysis toolkit for en face retinal vessels appearing in infrared reflectance scanning laser ophthalmoscopy (SLO) images.
Methods SLOctolyzer includes two main modules: segmentation and measurement. The segmentation module use deep learning methods to delineate retinal anatomy, while the measurement module quantifies key retinal vascular features such as vessel complexity, density, tortuosity, and calibre. We evaluate the segmentation module using unseen data and measure its reproducibility.
Results SLOctolyzer’s segmentation module performed well against unseen internal test data (Dice for all-vessels, 0.9097; arteries, 0.8376; veins, 0.8525; optic disc, 0.9430; fovea, 0.8837). External validation against severe retinal pathology showed decreased performance (Dice for arteries, 0.7180; veins, 0.7470; optic disc, 0.9032). SLOctolyzer had good reproducibility (mean difference for fractal dimension, -0.0007; vessel density, -0.0003; vessel calibre, -0.3154 $\mu$m; tortuosity density, 0.0013). SLOctolyzer can process a macula-centred SLO image in under 20 seconds and a disc-centred SLO image in under 30 seconds using a standard laptop CPU.
Conclusions To our knowledge, SLOctolyzer is the first open-source tool to convert raw SLO images into reproducible and clinically meaningful retinal vascular parameters. It requires no specialist knowledge or proprietary software, and allows manual correction of segmentations and re-computing of vascular metrics. SLOctolyzer is freely available at https://github.com/jaburke166/SLOctolyzer.
Translational Relevance SLO images are captured simultaneous to optical coherence tomography (OCT), and we believe our software will be useful for extracting retinal vascular measurements from large OCT image sets and linking them to ocular or systemic diseases.
1 Introduction
The retina is a highly vascularised light-sensitive tissue at the back of the eye. The retinal microvasculature branches from the optic nerve head along the inner surface of the retina and is commonly imaged using colour fundus photography, which typically covers a 35–50 degree field of view around the posterior pole.
Confocal near infrared reflectance scanning laser ophthalmoscopy (SLO) utilise long wavelengths of around 820nm and also captures an en face view of the superficial retinal vessels 1. SLO images are typically captured simultaneous to optical coherence tomography (OCT) scans of the inner retinal layers (supplementary Fig. S1) acting as a localiser to position the OCT scan at the back of the eye accurately and typically images a restricted field of view of 30 degrees (approximately 9 mm2). Additionally, confocal imaging leverages laser-point raster scanning to prevent stray light interference during acquisition 2, thus providing higher contrast visualisation of the en face retinal vasculature over regular colour fundus images. Moreover, the interferometry of OCT imaging enables an exact conversion from pixel to physical space, a crucial step in forming clinically meaningful datasets of features of the en face retinal vessels 3. In colour fundus images, heuristic approaches are commonly used to obtain physical measurements, such as optic disc area normalisation 4. However, despite the detail SLO images capture they are often discarded due to the primary aim of OCT imaging to visualise the cross-sectional retinal layers.
This improved visualisation, increased accuracy in measurement and growing abundance from OCT capture make the SLO imaging modality an exciting frontier of en face retinal vessel analysis. Manual delineation of en face retinal vessels is prohibitively time-consuming, labour-intensive, and prone to inaccuracies. Consequently, extensive research has focused on segmenting retinal vessels, primarily in the colour fundus modality 5, with less emphasis on SLO images.
At the time of writing, very few methods have been reported for retinal vessel analysis for SLO images 6; 7; 8; 9; 10, and none for other major landmarks on the SLO image, the fovea and optic disc. Additionally, none of these methods have been made available to the research community, which has the potential to impact the standardisation of SLO-derived retinal vascular parameters.
Equipping automatic imaging methods with the tools for measuring downstream features for research and clinical use is an important element of any research pipeline 11; 12; 13; 14; 15. This allows the general researcher to use these methods in an end-to-end manner, and analyse their own retinal images without specialist training or a background in image processing. Unfortunately, there is a lack of similar pipelines for SLO images.
Accordingly, we aimed to develop and release a fully automatic analysis toolkit, SLOctolyzer, for segmentation and subsequent measurement of retinal vessels as seen on SLO images. Here, we introduce and quantitatively validate the segmentation models which detect the binary vessels, and classification of arteries and veins, as well as detection of the optic disc and fovea.
We highlight several contributions that SLOctolyzer provides to the research community:
  • SLOctolyzer is the first open-source pipeline for converting raw SLO images into clinically meaningful, retinal vascular parameters.
  • SLOctolyzer’s measurement module computes several summary measures of the retinal vessels, such as fractal dimension, vessel density, tortuosity density, vessel calibre and central retinal artery/vein equivalents.
  • SLOctolyzer works on macula- or optic disc-centred images which usually accompany OCT capture.
  • SLOctolyzer supports manual annotation to correct erroneous segmentations.
  • SLOctolyzer is suitable for batch processing large amounts of SLO images, collating results for ease of segmentation quality inspection and downstream statistical analysis.
2 Methods
Fig. 1 summarises the key elements of SLOctolyzer’s pipeline, which contains a segmentation module and a measurement module (of which the latter is based on features and code produced by Automorph 12). We have released SLOctolyzer in a frictionless manner so that it can be used without author permissions, proprietary software, or specialist training. SLOctolyzer is freely available here: https://github.com/jaburke166/SLOctolyzer.
Figure 1: Summary of SLOctolyzer’s analysis pipeline with a segmentation module (a) for binary vessel/artery/vein segmentation and fovea/optic disc detection, and a measurement module (b) for measuring retinal vascular parameters, with additional functionality for metadata extraction and manual annotation.
2.1 Data
Data used to build the segmentation module include RAVIR 10, a publicly available dataset of pathological SLO images with the arteries and veins segmented, and three additional cohorts related to systemic health: i-Test 16, a cohort of pregnant women at late gestation who are experiencing either a normative, pre-eclamptic or small baby pregnancy; PREVENT-Dementia 17; 18, a cohort of mid-life individuals, half of whom are at genetic risk of developing later life dementia; FutureMS 19; 20, a cohort of individuals with newly diagnosed relapsing-remitting multiple sclerosis (MS). All studies and cohorts adhered to the Declaration of Helsinki and received relevant ethical approval and informed consent from all subjects was obtained in all cases from the host institution.
Table 1 summarises image characteristics of each cohort. All SLO images were captured using a spectral domain Heidelberg SPECTRALIS HRA+OCT imaging device (Heidelberg Engineering, Heidelberg, Germany). All images were either centred on the macula or the optic disc. SLO images from Heidelberg HRA+OCT imaging devices are grayscale, capturing a 30 degrees field of view (approximately 9 mm2) using a laser light with wavelength 820nm. All cohorts used a HRA+OCT SPECTRALIS Standard Module except for the i-Test study, which used a portable version, the SPECTRALIS FLEX, for 24 participants out of the 93 available. Dependent upon settings prior to acquisition, SLO images from Heidelberg devices have either a pixel resolution of 768 $\times$ 768 or 1536 $\times$ 1536 (pixel height $\times$ width).
Study Participants Images Right eyes Retinal pathology HRA+OCT Module Image resolution Location
RAVIR 10 23 23 14 Yes Standard 768 Disc
Healthy 21 7 7 7 No Standard 1536 Disc
PREVENT 17; 18 144 285 142 No Standard 1536 Disc
i-Test 16 93 186 93 No Standard & FLEX 768 Macula
FutureMS 19; 20 15 15 9 No Standard 1536 Disc
Table 1: Image characteristics of the five cohorts used to build SLOcolyzer’s segmentation module. Image resolution is in pixels (for both lateral and axial directions), location refers to the centring of the scan, i.e. if it’s macula-centred disc-centred. HRA, Heidelberg retina angiography; OCT, optical coherence tomography.
SLOctolyzer’s segmentation module contains three segmentation models, one for binary vessel detection, one for fovea detection and another for artery-vein and optic disc detection. Table 2 outlines which cohorts and how many eyes/images of the cohorts each model used for training and evaluation. For all studies except for FutureMS, we selected all available participants for modelling. For FutureMS, a subset of 15 participants from the wider cohort 20 were selected. These images were chosen based on image features relevant for vessel segmentation, such as blur, illumination, contrast, and abnormal features like vessel tortuosity or optic nerve head atrophy (supplementary Fig. S2). Similarly, for the i-Test cohort, 15 participants were selected from the wider cohort for building the artery-vein-optic disc (AVOD) model.
Model Cohort (eyes/images) Total
RAVIR Healthy PREVENT i-Test FutureMS
Vessel 23 7 30
Fovea 23 7 285 186 15 516
AVOD 15 15 30
Table 2: Number of eyes (images) used for each segmentation model stratified by cohort. AVOD, artery-vein-optic disc.
2.2 Ground truth labels
2.2.1 Binary vessel detection
A total of 30 SLO images (RAVIR & Healthy) were used to train and evaluate the segmentation model for binary vessel detection. 23 of these were sourced from the RAVIR dataset, of which manual pixel-level annotation of the arteries and veins were done by the same, experienced retinal image grader 10. Segmentation over the optic nerve head was only performed if artery and vein classes were able to be resolved. Vessel labels were defined as both the artery and vein class combined. The remaining seven SLO images were made available from healthy eyes and were manually segmented by an experienced image analyst (author S.G.) using ITK-Snap 22, which supports pixel-level annotation.
2.2.2 Fovea detection
Fovea detection was performed by the same experienced grader (author J.B.) and used all 516 SLO images made available.
For the i-Test cohort, access to both the SLO and OCT was possible, which allowed the fovea coordinate on the SLO to be cross-referenced using it’s location on the corresponding fovea-centred OCT B-scan. The fovea location on each OCT B-scan was detected using Choroidalyzer 15, a fully automatic toolkit for choroid analysis and fovea detection in OCT images. For the remaining images, only the SLO was available and the fovea coordinate was selected manually, using an in-house GUI written in Python (version 3.11.6).
We assessed intra-rater variability by one user (author J.B.) through re-selecting the fovea pixel coordinate for all 516 images, two months after the initial manual grading, and the results of which can be found in supplementary Table S1. Repeatability against the initial i-Test fovea coordinates acted as comparison to ground truth, and the remaining images were used to assess intra-rater variability. Intra-rater variability had an average intra-class correlation (ICC(3,1)) of 0.9934 for both $x$ and $y$ coordinates, and comparison to ground truth scored an average ICC of 0.8206 across both $x$ and $y$ coordinates.
The final fovea coordinate used for modelling was computed as the average of the two pixel coordinates selected.
2.2.3 Artery-vein-optic disc detection
Three experienced image analysts (authors J.B., S.G., A.T.) were given 14 SLO images to manually annotate the arteries, veins, and optic disc. For each SLO image, a binary vessel map was generated using the binary vessel detection model. Each grader then used ITK-Snap 22 for anatomical segmentation of arteries, veins and the optic disc. See supplementary Fig. S3 and S4 for the vessel labelling rules and segmentation protocol each rater followed for consistent and comparative segmentation.
Each set of 14 images had two images with distinct overlap between pairwise graders. Thus, there were a total of 30 unique SLO images used for modelling: 15 from the i-Test cohort and 15 from the FutureMS cohort. This overlap allowed for a total of six SLO images which could be used for comparison. We calculated inter-rater agreement for all-vessel, artery, vein, and optic disc segmentation using the Dice coefficient, whose results can be found in supplementary Table S2. All values were found to be greater than 0.93, suggesting an excellent degree of consistency between graders.
We also had a clinical ophthalmologist (I.M.) qualitatively rate the artery and vein segmentations for all 30 SLO images. For each SLO image, we asked I.M. to rate its image quality and the segmentation quality of the artery and vein annotation using a 5-point ordinal scale between -2 (very bad) and 2 (very good). The results of this rating can be found in supplementary Table S3. 63% (19/30) were rated as ‘very good’, with the remaining annotations graded as ‘good’, except for only one which was rated as ‘okay’. This example can be seen in supplementary Fig. S5, which had two, small miss-labelled vessels. Importantly, there were no ‘bad’ or ‘very bad’ ratings for these manual annotations.
2.2.4 The RAVIR dataset
The RAVIR dataset 10 was intended to be used for training the artery-vein-optic disc (AVOD) model, as it contains suitable examples of severe retinal pathology. However, before modelling we reviewed the segmentation labels and observed inconsistent artery and vein classification, an example of which can be seen in supplementary Fig. S6. Therefore, we decided to use the RAVIR dataset as an external test set for the AVOD model, after correcting for any misclassified veins or arteries. This external test set presents a significant challenge to the AVOD model as it contains examples of severe retinal pathology.
Before evaluating the AVOD model, the corrections to the RAVIR dataset were reviewed by a clinical ophthalmologist (I.M.) against the original labels in a masked, randomised fashion. For each image, I.M. was presented with two segmentation labels without revealing their source. He was then asked to rate the SLO image quality and each artery-vein classification using a 5-point ordinal scale from -2 (very bad) to 2 (very good), and select which of the labels he preferred, with options for none or both. The results are shown in supplementary Table S4, and the ophthalmologist preferred over 64% (14/23) of the corrected labels, and the labels for the remaining images were equally of good quality, according to I.M. Additionally, we segmented the optic disc for this external test set for comparison with our optic disc segmentation.
2.3 Segmentation module
During training, we applied the same data augmentations randomly to each model’s training data. This included random horizontal and vertical flips ($P=0.5$), affine transformations to allow random rotation and scaling, uniformly sampled in $[0,90]$ and $[0.4,1.6]$, respectively ($P=1/3$). We also applied a Gaussian blur with random window size and standard deviation uniformly sampled in $[1,20]$ and $[0.1,4.0]$, respectively ($P=1/4$). Finally, we applied random brightness and contrast adjustments, altering the brightness and contrast using uniformly sampled scale factors $[0.8,1.2]$ and $[0.75,1.25]$, respectively.
We used python 3.12, PyTorch 2.0, Segmentation Models PyTorch 23 and MONAI 24.
2.3.1 Binary vessel detection
The model and training pipeline has been discussed previously 25. Briefly, the model was a custom U-Net deep learning architecture trained from scratch on a total of 30 SLO images. The dataset was split randomly into training (n=24), validation (n=2) and test (n=4) sets. Offline data augmentation randomly cropped 20 patches of size 320 $\times$ 240 to artificially increase the training set, reduce computational load and accelerate the training rate.
The model was trained to minimise the Dice Focal loss 24, using the Adam 26 optimiser (with default optimiser hyperparameters) for 600 epochs using a batch size of 20. The initial learning rate was set as 1e-3 until epoch 300, at which point the learning rate was reduced to 1e-4 for the remainder of training.
2.3.2 Fovea detection
Fovea segmentation is a significantly easier task than vessel detection or artery-vein classification, so our model used a pre-trained 23, lightweight U-Net deep learning architecture with a MobileNetV3 27 backbone, previously trained using ImageNet 28. This model used all available images from all cohorts, comprising 516 SLO images.
The dataset was split into training ($n=359$), validation ($n=104$) and test sets ($n=54$) at the participant-level, so there was no overlap of participants between sets. All images were resized to a common pixel resolution of $768\times 768$ for training. To facilitate learning through segmentation, training labels were circular binary masks whose centre was the fovea coordinate with radius of 60 pixels.
The model was trained to minimise the binary cross entropy loss using the AdamW 29 optimiser for 200 epochs using a batch size of 8. During training, we applied two sets of random augmentations per image, per epoch, artificially doubling the batch size. An initial learning rate of 5e-4 was set, and after a 10-epoch linear warm up to a learning rate of 5e-3, it decayed following a cosine relationship for 90 epochs, with the learning rate set at 5e-4 for the remaining 100 epochs.
2.3.3 Artery-vein-optic disc detection
The AVOD model used 30 SLO images from two cohorts related to pregnancy (i-Test, 15 images) and multiple sclerosis (FutureMS, 15 images). Given the number of segmentation tasks and dataset size, we similarly opted for fine-tuning a pre-trained U-Net deep learning model, this time using a deep ResNet101 30 backbone, previously trained on ImageNet 28. The dataset was split into training (n=20), validation (n=4) and test (n=6) sets, with equal proportion of cohorts (and thus macula- and disc-centred images) in each set. All images were resized to a common pixel resolution of 768 $\times$ 768 for training.
The model was trained to minimise the Dice Focal loss 24, using the Adam 26 optimiser (with default optimiser hyperparameters) for 300 epochs using a batch size of 4, and a learning rate of 5e-4.
2.4 Measure module
SLOctolyzer’s second module is a suite of tools for measuring retinal vascular parameters, and are based on the features measured when processing colour fundus images using Automorph 12.
We use the whole image as the region of interest for feature measurement, regardless of the location of the SLO image. Additionally, for disc-centred SLO images, we define two concentric rings centred at the optic disc, zones B and C 21. Zone B takes an annulus from one-half diameter away from the optic disc margin to a full diameter away (0.5D - 1D), and zone C measures from the optic disc margin to two diameters away (0D - 2D). See Fig. 1(B) for a visualisation of these zones.
Across the whole image, we measure global features of vessel complexity, density and calibre of the vasculature, using fractal dimension (Minkowski-Bouligand dimension) 31, vessel density (ratio of vessel pixels to all pixels) and global vessel calibre (ratio of vessel pixels to skeletonised vessel pixels. These measurements are not taken for localised zones B and C. For every region of interest (whole image and zones), we measure tortuosity density 32 and local vessel calibre across individual vessel segments, whose endpoints are defined by arteriovenous crossing and bifurcations. Global and local vessel calibre are very similar measures, but the latter provides a more granular approach to computing calibre from individual vessel segments — see supplementary Fig. S7 and supporting text for more details.
All features are measured across artery, vein and binary vessel maps. Finally, for artery and vein vessel maps, we measure central retinal artery and vein equivalents (CRAE/CRVE) using the Knudston approach 33, as well as the arteriole-to-venule (AVR) ratio. While these metrics are typically only measured for zones B and C in disc-centred images, they are also computed across the whole image and thus for macula-centred images too.
2.5 SLOctolyzer’s pipeline
Fig. 1 describes the analysis pipeline for SLOctolyzer. During segmentation, the image is resized to a common pixel resolution of 768 $\times$ 768 as the segmentation module was trained on this common pixel resolution. This prevents any segmentation error in larger images due to the central artery/vein light reflex 25, which appears as a bright strip in the centre of a vessel, potentially leading to a larger vessel being perceived as two, smaller and parallel ones. Upon resizing, full segmentation of the vasculature, optic disc and fovea is performed. After thresholding with a value of 0.5 to produce binary vessel masks, post-processing is applied to remove small false positive regions, and disconnected vessels are joined to their neighbouring vessels to remove small gaps. The fovea coordinate is defined as the centroid of the fovea binary mask. If the SLO image is disc-centred, the optic disc binary mask is modelled as an ellipse, and it’s diameter is defined as the average of the major and minor axis lengths.
After segmentation, if the SLO image is macula-centred, measurements on the vessel/artery/vein maps are taken across the whole image. If the SLO image is disc-centred, the measurements are taken across the whole image and for zones B and C, which are defined using the computed optic disc diameter. These measurements will be saved out per file, alongside key metadata either stored in the input file, or inferred during analysis (laterality, location, fovea centre, optic disc diameter and centre). Moreover, segmentation masks will be saved out, as well as helpful visualisations to inspect segmentation quality. If running on a batch of images, a collated file will be saved out row-wise compiling all measurements and metadata from all image files analysed. Finally, a folder of visualisations, superimposing the segmentations onto the SLO image will be saved out per file which will aid segmentation quality checking.
While we do not have an automatic way to assess segmentation quality, an additional feature of SLOctolyzer is its ability to correct manual segmentations. Given a saved out segmentation mask of the arteries, veins and optic disc, the user may correct any erroneous areas using ITK-Snap 22 and save the corrected version in the same folder as the original mask. Upon rerunning the pipeline again, it will use this corrected annotation and re-compute the measurements for that image. See supplementary Fig. S8 and Fig. S9 with supporting text for more information on SLOctolyzer’s setup, interface and output.
2.6 Statistical analysis
Each segmentation model, after model selection based on the performance against their respective validation set, was evaluated on their respective, internal test set. Additionally, for the AVOD model, we used the RAVIR dataset 10 as an external test to measure the segmentation module’s robustness to severe retinal pathology.
We used standard segmentation metrics such as area under the receiver operating characteristic curve (AUC) and the dice similarity coefficient (at a standard threshold of 0.5) to evaluate segmentation performance. We also used mean absolute error (MAE) to measure performance on downstream, retinal vascular parameters fractal dimension and local vessel calibre (in microns, $\mu$m, where pixel length-scales were available and pixels, px, otherwise). We used MAE to measure the error in predicted fovea pixel coordinate (along both axes) and optic disc area.
For internal test set evaluation, the whole image was used as the region of interest due to the low sample size of disc-centred images in each internal test set. For external test set evaluation, all regions of interest (whole image, zone B and zone C) were evaluated, as RAVIR contains only disc-centred images.
We also measured the reproducibility of SLOctolyzer on downstream retinal vascular parameters, using the first 60 participants (120 eyes) of the i-Test cohort (whose data was available at the time of analysis). SLO images were captured simultaneous to macula-centred, OCT volume capture when enhanced depth imaging mode was toggled on and off, creating a repeated, albeit unregistered, pair of SLO images per eye. We measured all retinal vascular parameters previously mentioned across the whole image, and report mean absolute error and Pearson, Spearman and intra-class correlation (ICC(3,1)) coefficients for population-based reproducibility analysis. We also report Bland-Altman plots 34 to assess the distribution of residuals.
Additionally, we report the reproducibility of SLOctolyzer at the eye-level using a measure of individual-level measurement noise $\lambda$ 35. We express the variability of each measured feature within an eye in units of the feature’s overall population variability. $\lambda$ is computed per feature, per eye, as the ratio between the standard deviation of within-eye measurements and the standard deviation of between-eye measurements, across the population. For convenience, $\lambda$ is presented as a percentage, with 0 as the optimal value. Note that the population are all white women with a mean (SD) age of 34.7 (5.2) years, at mean (SD) gestation of 35.6 (3.4) weeks, and thus we do not expect there to be huge variation across the population.
Finally, we measured the execution time of the SLOctolyzer’s segmentation module and the whole pipeline on macula- and disc-centred images of different pixel resolutions. We ran each experiment 100 times and report the mean and standard deviation execution time in seconds. Timed experiments were run on a standard Windows laptop with a four year old i5 CPU with 16GB of RAM, but no GPU.
3 Results
Metric Vessel Fovea Artery Vein Optic disc
AUC Dice AUC Dice AUC Dice AUC Dice AUC Dice
Segmentation 0.9922 0.9097 0.9975 0.8837 0.9860 0.8376 0.9857 0.8525 0.9976 0.9430
FD Calibre [px] Cx [px] Cy [px] FD Calibre [px] FD Calibre [px] Area [px2]
Feature 0.0130 0.1518 6.2778 4.7778 0.0178 0.2081 0.0133 0.1539 638.1667
Table 3: Segmentation results for each segmentation task (top). Mean absolute error across relevant measurements for each segmentation task (bottom). Average local calibre was used here (in pixels). FD, fractal dimension. Cx is the mean absolute error in the horizontal component of the centroid for the fovea (and similarly for Cy.
Figure 2: Segmentation masks for each segmentation model on two examples randomly selected from their test sets.
3.1 Evaluation
Segmentation results are presented in Table 3 for all three models for their respective internal test sets. All models had a strong AUC score (all greater than 0.98), suggesting the raw probability maps were well calibrated in being able to score highly the relevant pixels related to each of their segmentation tasks, with the fovea and optic disc detection performing best, which was to be expected since these are easier segmentation tasks than vessel detection. Upon thresholding, both binary vessel and AVOD models scored high dice scores for all-vessel detection (binary vessel model: 0.9097, AVOD model: 0.8570), and the AVOD model scored slightly lower dice scores for artery and vein classification (artery, 0.8376; vein, 0.8525). This was unsurprising, since distinguishing between artery and vein is significantly more challenging than differentiating retinal vasculature from retinal tissue. Fig. 2 shows example segmentation outputs from the internal test sets.
The MAE for fovea coordinate detection was 6.3 and 4.8 pixels for the horizontal and vertical direction (Table 3), thus a MAE of 7.9 pixels diagonally. Images had a pixel resolution of 768 $\times$ 768, covering approximately a 9 mm2 field of view, which makes their conversion factor approximately 11.71 microns-per-pixel in both axes. Thus, this diagonal MAE corresponds to 92.7 microns, which is well within the average size of the foveola of 350 microns 36.
Similarly, an MAE of 638 square pixels for optic disc area corresponds to approximately an MAE of 0.08 mm2, which is not clinically significant given the wide distribution of optic disc sizes 37, the observed change across the myopic spectrum 38 and even differences across different devices 39. MAE for local vessel calibre was around 0.2 pixels, which corresponds to approximately 2.32 microns, and is similarly an insignificant error given population distributions for arterioles and venules using confocal SLO imaging 40.
The results of our segmentation module evaluated against the external test set is shown in Table 4. The RAVIR dataset 10 presents a significant challenge due to retinal pathology heavily featuring in the 23 SLO images. The model performs worse compared with the unseen test data which relate to systemic health (dice coefficient averaged across zones: artery, 0.7228; vein, 0.7598; optic disc, 0.9032), but quantitatively performs well for those images with mild pathology. Given the segmentation models were trained on images with grossly normal retinae related to systemic health changes, it is perhaps no surprise that the performance of the model decreases as the extent of retinal pathology increases. Fig. 3 compares the ground truth labels (middle row) with our ones (bottom row), for SLO images of increasing retinal pathology from left to right. While the major vessels are classified correctly for the majority of cases, as the pathology becomes more severe we see smaller and more tortuous vessels ignored, and segmentations of the major vessels misclassified.
Region of interest Artery Vein Optic disc
AUC Dice FD Width [px] AUC Dice FD Width [px] AUC Dice Area [px2]
Whole 0.9421 0.7180 0.0444 0.6869 0.9558 0.7470 0.0737 0.9791 0.9720 0.9032 788.5000
Zone B 0.9922 0.7300 0.0379 1.0125 0.9935 0.7726 0.0597 1.1823 - - -
Zone C 0.9643 0.7203 0.0440 0.7854 0.9702 0.7598 0.0543 1.1479 - - -
Table 4: Segmentation and feature measurement mean absolute errors of the artery-vein-optic disc (AVOD) segmentation model against the RAVIR external test set. Average local calibre was used here (in pixels). AUC, area under the receiver operating characteristic curve; FD, fractal dimension (non-dimensional).
Figure 3: Visual comparison between ground truth and predicted segmentations for the external test set, RAVIR, for artery-vein-optic disc detection. The images were selected to show progressing eye pathology from left to right.
3.2 Reproducibility
Table 5 present correlation values for all retinal vascular parameters for the binary vessel segmentation model (all-vessel), and the artery-vein-optic disc model. All correlations showed strong or excellent agreement among features, with fractal dimension and vessel density scoring the highest. Across the cohort, veins appeared as the poorest performing, but not by any significant margin.
Metric All-vessel Artery Vein
MAE Pearson/Spearman ICC(3,1) MAE Pearson/Spearman ICC(3,1) MAE Pearson/Spearman ICC(3,1)
Fractal dimension 0.0077 0.9326/0.9192 0.965 (0.95, 0.98) 0.0048 0.9485/0.9385 0.9734 (0.96, 0.98) 0.0070 0.8940/0.8860 0.9423 (0.92, 0.96)
Vessel density 0.0013 0.9493/0.9513 0.974 (0.96, 0.98) 0.0022 0.9393/0.9371 0.9686 (0.95, 0.98) 0.0014 0.9370/0.9218 0.9665 (0.95, 0.98)
Global calibre [$\mu$m] 1.5811 0.8771/0.8861 0.9293 (0.9, 0.95) 1.5995 0.8425/0.8331 0.9125 (0.87, 0.94) 1.6063 0.8654/0.8692 0.9276 (0.9, 0.95)
Local calibre [$\mu$m] 1.5049 0.8707/0.8815 0.9259 (0.89, 0.95) 1.4654 0.8489/0.8525 0.9165 (0.88, 0.94) 1.4105 0.8695/0.8711 0.9299 (0.9, 0.95)
Tortuosity density 0.0324 0.7431/0.7460 0.8526 (0.79, 0.9) 0.0147 0.7962/0.7977 0.886 (0.84, 0.92) 0.0241 0.7039/0.6825 0.8262 (0.75, 0.88)
AVR/CRAE/CRVE [$\mu$m] 0.0283 0.7743/0.7931 0.8728 (0.82, 0.91) 6.2834 0.8456/0.8554 0.9151 (0.88, 0.94) 10.1025 0.7940/0.8009 0.8852 (0.83, 0.92)
Table 5: Population-based reproducibility performance of SLOctolyzer, reporting mean absolute error (MAE), Pearson, Spearman and intra-class correlations (ICC(3,1)). All Pearson and Spearman correlations were statistically significant with P-values $P<0.0001$. For readability, AVR/CRAE/CRVE were combined into one row — AVR is dimensionless, while CRAE and CRVE are in microns and only report for arteries and veins, respectively.
Fig. 4 present Bland-Altman plots of the residuals between paired SLO images. We found that SLOctolyzer was highly reproducible, with the distribution of residuals according to most metrics centred around 0, showing no apparent trend. For fractal dimension, there was a mean difference (MD) of -0.0007 with limits of agreement (LoA) of [-0.0178, 0.0164]; for vessel density there was a MD of -0.0003 and LoA of [-0.0052, 0.0046]; for global vessel calibre there was a MD of -0.3154 $\mu$m and LoA of [-4.9804, 4.3496]; for local vessel calibre there was a MD of -0.3208 $\mu$m and LoA of [-4.5777, 3.9361]; for tortuosity density there was a MD of -0.0013 and LoA of [-0.0624, 0.0650]; for CRAE there was a MD of 0.9901 $\mu$m and LoA of [-17.4345, 15.4543]; for CRVE there was a MD of -1.8782 $\mu$m and LoA of [-30.3251, 26.5687]; finally, for AVR there was a MD of 0.0011 and LoA of [-0.0738, 0.0760].
Figure 4: Bland-Altman analysis of the residuals from assessing the reproducibility of the segmentation models for all features outputted by SLOctolyzer. Residuals for arteries (red), veins (blue) and all-vessel (orange) are plotted together.
Fig. 5 reports the eye-level reproducibility of each of the features measured by SLOctolyzer. For fractal dimension, vessel density and global/local calibre, the upper, or third quartile is below 25% across all vessel types. Thus, the range of measurement error for these features in the majority of the paired, but unregistered, eyes were below 25% of the overall populations’ variability. Given the homogeneous nature of the population’s demographics, this is very reasonable. Moreover, vessel density has the smallest interquartile range among features (mean all-vessels, $\lambda=11.7\%$; arteries, $\lambda=12.1\%$; veins, $\lambda=13.3\%$), suggesting this feature, closely followed by fractal dimension, had the least measurement noise. Tortuosity density and AVR/CRAE/CRVE appear the most sensitive features, with veins as the most variable relative to arteries and all-vessels.
The eyes with the worst reproducibility for vessel calibre and tortuosity density are shown to the right of the box plots, with dashed lines linking the major outlier scatter-points to the corresponding eye. For all calibre metrics, the worst metric came from the same eye. This was due to an arteriovenous crossing near the edge of the image, where a major artery and vessel overlapped, and SLOctolyzer classified artery and vein differently (red arrows). This led to a large absolute difference of 87.5 microns in CRVE. For tortuosity density, the vessel maps do not look qualitatively different, but the absolute difference was 0.096. We believe that this is due to tortuosity density’s sensitivity to arteriovenous crossings which disconnect the respective artery and vein vessel maps. Star-shaped scatter-points represent second to largest outliers for vessel calibre and tortuosity density and are shown in supplementary Fig. S10, and are similar due to arteriovenous crossings or vessel disconnectedness.
Figure 5: Reproducibility at the eye-level of each of the features outputted by SLOctolyzer. Box-plot distributions are shown for each feature, with all-vessels in orange, arteries in red and veins in blue. Note that the AVR/CRAE/CRVE features are presented together as CRVE/CRAE are measured across arteries (red) and veins (blue) individually, while AVR is a combination of the two, thus represented in the all-vessel class (orange). A horizontal, dashed line at the $\lambda$=25mark is shown as a visual aid. We also show the eyes (with segmentations overlaid) which had the greatest reproducibility measurement error for large-vessel calibre and vein tortuosity density (circular scatter-points). Eyes corresponding to the star-shaped scatter-points are shown in supplementary Fig. S10.
3.3 Execution time
Table 6 shows the execution time of SLOctolyzer’s segmentation module and entire pipeline. Because the segmentation module resizes all images to a common resolution of 768 $\times$ 768, the segmentation execution time across all image types and resolutions are approximately equal. For binary vessel segmentation, it takes approximately 4.5 s/image, for the lightweight fovea detection model it takes approximately 0.6 s/image, and for artery-vein-optic disc detection it takes approximately 3.0 s/image. Thus, SLOctolyzer’s segmentation module using a standard laptop CPU takes approximately 11.1 $\pm$ 0.5 s/image. Measuring the execution time of the entire pipeline on a standard laptop CPU took 18.9 $\pm$ 1.0s for a macula-centred SLO image at 768 $\times$ 768 resolution, 28.6 $\pm$ 1.0s for a disc-centred SLO image at 768 $\times$ 768 resolution and 111.0 $\pm$ 3.5s for a disc-centred SLO image at 1536 $\times$ 1536 resolution. Disc-centred SLO images are more time intensive for feature measurement as measurements are made across the whole image, and for zones B and C across the vessel, artery and vein segmentation maps separately.
Image type Segmentation module Whole pipeline
Binary vessel Fovea AVOD
Macula-centred ($768\times 768$) 4.8 (0.407) 0.6 (0.067) 2.9 (0.408) 18.9 (0.984)
Disc-centred ($768\times 768$) 4.3 (0.147) 0.6 (0.047) 2.9 (0.440) 28.6 (1.040)
Disc-centred ($1536\times 1536$) 4.2 (0.191) 0.6 (0.051) 2.7 (0.133) 111.0 (3.540)
Table 6: Average (standard deviation) execution time in seconds of the segmentation module and SLOctolyzer’s entire pipeline, for SLO images of different location and size.
4 Discussion
We have developed a fully automatic method, SLOctolyzer, for processing macula-centred and disc-centred SLO images which typically accompany OCT data but are often discarded in research. SLOctolyzer converts a raw SLO image into clinically relevant retinal vascular parameters, including fractal dimension, tortuosity, and vessel calibre.
SLOctolyzer performed well on unseen data, but we observed a performance drop against an external test set featuring severe retinal pathology. SLOctolyzer was designed for modelling the retinal vasculature during systemic health change, so there was understandably a noticeable decrease in performance as the subjective extent of pathology increased. However, we believe SLOctolyzer’s additional functionality to correct any annotations and re-compute measurements will prove very useful for cases of severe retinal pathology. Regardless, we envision SLOctolyzer being a very useful tool for SLO image analysis, particularly as the nascent field of oculomics increases rapidly 41, since the populations being studied often have grossly normal retinae.
SLOctolyzer’s segmentation module had good reproducibility, with the residuals across all metrics for vessels, arteries and veins exhibiting no pattern and being centred around 0. In particular, the residuals for measurements of global and local calibre (in microns) appear to show proportionately small error in comparison to what is generally considered as clinically significant. Meta-analyses have found vessel calibre difference of 10–15 microns between mid- and old-age, with 20 micron increase in venule calibre associated with increased risk of stroke by 1.15 42.
The distributions of measurement error at the eye-level for fractal dimension, vessel density and local/global calibre were found to be reasonable given the homogeneous nature of the i-Test cohort studied for reproducibility — with tortuosity density and large-vessel calibre having higher variability. We observed that small differences in the segmentation maps can have an inordinate impact on tortuosity density (supplementary Fig. S10(A)). This highlights the apparent sensitivity of tortuosity, regardless of whether the segmentation looks qualitatively different. While segmentation error contributes to this error, this observation adds to a potential problem surrounding standardisation of retinal vascular parameters 43. For large-vessel calibre, the largest outliers for large-vessel calibre came from the same eye, which showed a major arteriovenous crossings at the edge of the image causing a major artery-vein miss-classification. Moreover, large-vessel calibre metrics like AVR/CRAE/CRVE are typically reserved for disc-centred SLO images as these show the major vessels in full. Thus, the unregistered nature of the macula-centred SLO pairs, with major vessels appearing toward the edge of the image, also contributed to the measurement error observed (supplementary Fig. S10(B)). We hope in the future we may also assess SLOctolyzer’s reproducibility for paired, disc-centred SLO images.
On a standard laptop CPU, SLOctolyzer can segment an SLO image in around 11 seconds. For a standard 768 $\times$ 768 resolution image, the entire pipeline takes 19 seconds for macula-centred images and 29 seconds for disc-centred images. It can be used interactively or for batch processing (supplementary Fig. S8), converting raw SLO images into reproducible vascular metrics and outputting intermediary segmentation masks for quality inspection (supplementary Fig. S9).
There were some limitations associated with this work. Firstly, the SLO images used to build SLOctolyzer’s segmentation module were sourced from only one manufacturer, Heidelberg Engineering. Thus, we do not know the extent of performance from SLO images from other devices. Furthermore, most of the images we used were from individuals recruited to studies of systemic health and did not have obvious retinal pathology. We therefore expect the segmentation module to not be entirely robust to cases of severe retinal pathology, as seen by the performance against the RAVIR dataset (Table 4), which featured severe retinal pathology. However, SLOctolyzer’s additional functionality allows for segmentation error to be corrected manually.
A technical limitation of the work is that SLOctolyzer’s segmentation module is made up of three distinct models. The rationale for this was clear at the time of model construction: we could leverage an initial binary vessel segmentation model to help us manually annotate the arteries and veins for building an artery-vein detection model, a much more challenging task, with the fovea detection model utilising all available images due to the speed of foveal pit manual annotation. However, these models now provide a quick and consistent way to annotate large datasets which, after inspection and correction by expert graders, could be used to build a much more robust segmentation model.
Finally, our datasets were of macula-centred and disc-centred SLO images, and thus we do not know its performance at different locations on the retina.
We intend to improve SLOctolyzer’s segmentation module in future work by constructing a single model to perform all segmentation tasks and refine the dataset on which it will be trained and evaluated on. This will ideally include more diverse SLO images, such as those related to retinal pathology, capturing different regions of the retina, and from multiple imaging devices and manufacturers. Improving SLOctolyzer’s ability to generalise to other SLO imaging modalities, such as Optos ultra-wide field SLO 44, or Heidelberg’s more recent multicolour SLO 45, would further increase the scope of this open-source pipeline.
5 Conclusion
We have developed a fully automatic method for extracting useful information from SLO images, including fractal dimension, tortuosity, and vessel calibre. SLOctolyzer, the first open-source analysis toolkit for macula- and disc-centered SLO images, addresses the lack of open-source tools for this image modality. It leverages existing OCT imagesets to provide clinically meaningful retinal vascular parameters from the SLO images which are captured in parallel to OCT acquisition, but are often overlooked. Available on GitHub, SLOctolyzer requires no specialist training or proprietary software and supports batch processing and manual correction. We hope this tool will encourage the research community to reconsider the value of SLO images, facilitate collaboration and help to standardise retinal vascular parameters in SLO images and retinal imaging more generally.
Acknowledgements
J.B. was supported by the Medical Research Council (grant MR/N013166/1) as part of the Doctoral Training Programme in Precision Medicine at the Usher Institute, University of Edinburgh. For the purpose of open access, the authors have applied a creative commons attribution (CC BY) licence to any author accepted manuscript version arising.
6 Conflicts of Interest
The authors declare no conflicts of interest.
SUPPLEMENTARY MATERIALS
Simultaneous OCT + SLO capture during acquisition
Supplementary Fig. S1 shows a screenshot from the Heidelberg Eye Explorer (HEYEX) software (version 1.12.1.0) (Heidelberg Engineering, Heidelberg, Germany) of an OCT volume for an individual’s left eye. During OCT capture, the confocal SLO image (left) is used to reference the location of the B-scans (right) captured during acquisition. This is advantageous particularly for registering follow-up scans.
Figure S1: Screenshot from Heidelberg Eye Explorer (HEYEX) of an OCT volume capture during acquisition. (Left) the confocal SLO image with the acquisition location of the OCT volume overlaid in green. (Right) the corresponding fovea-centred OCT B-scan of the OCT volume.
SLOctolyzer’s training data for artery-vein-optic disc detection
Supplementary Fig. S2 shows some exemplary SLO images selected from the i-Test and FutureMS cohorts for building the artery-vein-optic disc segmentation model. The total set of 30 SLO images were selected to provide a variety of image-features during training and evaluation, such as blur, illumination, and contrast artefacts, as well as abnormal retinal features such as vessel tortuosity or optic nerve head atrophy.
Figure S2: Selection of the data used to build the artery-vein-optic disc segmentation model, highlight features of interest in the image which are advantageous for diverse training, such as blur (A), contrast artefacts (B,E,H), illuminations (D, F, G), vessel tortuosity (C) and optic nerve head atrophy (H).
Repeatability of fovea manual detection
As part of the manual detection of the fovea coordinate on SLO images, author J.B. manually selected the fovea coordinate for all cohorts but the i-Test one (as we could cross-reference the fovea on the SLO using the fovea-centred OCT B-scan). As a measure of rater variability, author J.B. re-selected the fovea coordinate for all 516 SLO images two months after the initial manual detection, and compared the $(x,y)$-positioning using mean absolute error and intra-class correlation. Like the training pipeline for the fovea detection model, fovea segmentation masks were generated for each batch of coordinates and compared using the area under the receiver operating characteristic curve (AUC) and dice coefficient. Evaluation was performed after all images were resized to a common pixel resolution of 768 $\times$ 768 for interpretable comparison.
The repeatability results are listed below in supplementary Table S1. Results for the i-Test cohort represent inter-observer variability (manual detection against accurate ground truth data), and results for the remaining four cohorts of data represent intra-observer variability. Intra-observer variability has an excellent average intra-class correlation of 0.9934 and a strong intra-class correlation for inter-observer variability of 0.8206.
AUC Dice ICC(3,1) MAE (px)
x y x y
Others 0.9640 0.9287 0.9998 0.9870 3.3921 3.0000
i-Test 0.9441 0.8892 0.7720 0.8692 4.7527 4.5161
Table S1: Repeatability of detecting the fovea on the i-Test dataset (which act as ground truth) and the remaining cohorts (which measure variability in the rater (J.B.). AUC, area under the receiver operating characteristic curve; ICC, intra-class correlation for single fixed raters; MAE, mean absolute error (in pixels)
Manual annotation of artery-vein-optic disc
The optic disc was classified as both the cup and rim together, and its margin was defined by the sharpest intensity transition between the darker rim and brighter retinal tissue constituting the retinal nerve fiber layer. For accurate artery-vein classification, the list below is a set of classification rules each grader followed:
  1. 1.
    Alternating rule: It’s common to see arteries and veins appear on the fundus in an alternative nature, for efficient oxygen transport 46.
  2. 2.
    Junction rule: It is common to observe arteriovenous crossings (a junction of four lanes) and same-class vessel bifurcations (a junction of three lanes). Given a crossing where there are more than three lanes, it’s likely to be an arteriovenous crossing. Otherwise, it is a same-class vessel bifurcation.
  3. 3.
    Thickness rule: Veins are typically thicker than arteries 40.
  4. 4.
    Brightness rule: Veins are typically darker than arteries 47.
Supplementary Fig. S3 shows an exemplary SLO image and annotations to describe each of the rules followed by the graders.
Figure S3: An exemplary SLO image with annotations to describe each of the four core rules for artery-vein classification.
A protocol was created to ensure consistent and comparable segmentations by each grader using ITK-Snap 22. After application of the binary vessel segmentation model to detect all the vessels in each SLO image, the core segmentation rules the graders followed are listed below:
  1. 1.
    Given an arteriovenous crossing, select the class which is most obviously overlaid on top the other class. If unsure, select the largest vessel.
  2. 2.
    Do not classify any vessels within the optic disc.
  3. 3.
    Complete any disconnected vessels.
  4. 4.
    Do not extend vessels which have visibly not been segmented.
  5. 5.
    Remove false positive pixels, i.e. retinal tissue which have been classed as vessel.
Supplementary Fig. S4 shows an exemplary SLO image with a binary vessel map overlaid using the binary vessel detection model, with annotations to highlight each of the four segmentation rules each grader followed.
Figure S4: An exemplary SLO image (A) with a raw binary vessel mask (B) superimposed in green with annotations to describe each segmentation rules followed during annotation.
Inter-rater agreement of graders for artery-vein-optic disc detection
Supplementary Table S2 shows the results of comparing the segmentations for each task (vessel/artery/vein/optic disc) across the three raters using the dice coefficient (ignoring background). All dice scores were greater than 0.93, suggesting strong agreement between graders.
Pairwise grader Vessel Artery Vein Optic disc
Grader 1 vs. Grader 2 0.9898 0.9597 0.9703 0.9442
Grader 1 vs. Grader 3 0.9906 0.9526 0.9602 0.9480
Grader 2 vs. Grader 3 0.9887 0.9468 0.9596 0.9379
Table S2: Inter-grader agreement between the three segmentation raters using the Dice similarity coefficient.
Qualitative evaluation of artery and vein ground truth labels
Supplementary Table S3 show the qualitative grading results from the clinical ophthalmologist (I.M.) for the manual annotations. All SLO images were graded as having ‘good quality’ or above, except for one, and there was no annotation which received a bad or worse rating. Across the image quality categories, a total of 19 SLO images were marked as ‘very good’, 10 as ‘good’ and 1 as ’okay’.
The SLO image and manual annotation which had the poorest score is shown in supplementary Fig. S5. The clinical ophthalmologist identified an arteriole and venule nasal to the optic disc as misclassified (red arrows). This error amounted to a drop in score from ‘good’ to ‘okay’, suggesting that all other images which rated as ‘good’ had significantly smaller miss-classifications, if any.
SLO Image quality Artery-vein rating Very good ($n=10$) VG: 7, G: 3, O: 0, B: 0, VB: 0 Good ($n=13$) VG: 7, G: 5, O: 1, B: 0, VB: 0 Okay ($n=7$) VG: 5, G: 2, O: 0, B: 0, VB: 0
Table S3: Qualitative adjudication of the manual annotations for artery and vein classification from a clinical ophthalmologist (I.M.). VG, very good; G, good; O, okay; B, bad; VB, very bad.
Figure S5: SLO image (A) and corresponding manual annotation superimposed (B) with poorest score from manual adjudication from clinical ophthalmologist (I.M.). The red arrows indicate where an arteriole/venule was miss-classified.
RAVIR segmentation label corrections
During the construction of the artery-vein-optic disc detection (AVOD) model, we intended on using the publicly available RAVIR dataset to further enhance and diversify our dataset of SLO images. However, during experimentation, we observed some inconsistent artery and vein classification.
One of our image graders (author J.B.) identified any major errors in the original segmentation labels and corrected them using ITK-Snap 22. Supplementary Fig. S6 shows an SLO image with the most significant error (A), the SLO image with the original labelling overlaid (B), and the SLO image with the corrected version overlaid (C).
After major corrections were made to all SLO images, we asked a clinical ophthalmologist (author I.M.) to rate the quality of the SLO image, as well as the segmentations before and after correction in a masked and randomised fashion using a 5-point ordinal scale from 2 (very good) to -2 (very bad). We also asked him to rate his preference, with options for both or neither available. Supplementary Table S4 shows the results from this adjudication, stratified by image quality.
Figure S6: SLO image (A), original manual annotation superimposed 10 (B) and corrected manual annotation (C). Here, many of the vessels in the macula were all assumed to be venules in the original annotation.
Image Quality Original 10 Corrected Both Very good ($N=6$) 0 3 3 Good ($N=4$) 0 2 2 Okay ($N=10$) 0 7 3 Bad ($N=3$) 0 2 1 Total ($N=23$) 0 14 9 Label Artery-Vein classification Original 10 VG: 6, G: 8, O: 6, B: 2, VB: 1 Corrected VG: 10, G: 11, O: 2, B: 0, VB: 0
Table S4: Qualitative adjudication between the original and corrected manual annotations for artery and vein classification in the RAVIR dataset 10, from a clinical ophthalmologist (I.M.). VG, very good; G, good; O, okay; B, bad; VB, very bad.
We chose to exclude the RAVIR dataset from training the artery-vein-optic disc model and used the corrected segmentation labels as an external test set for our model. We hypothesised this set would pose a significant challenge for our model, which was trained on images related to systemic health.
Global and local vessel calibre
Measurements of fractal dimension, vessel density and global vessel calibre are not taken across zones B and C because they capture a global measure of the vasculature. Instead, for zones B and C, we compute tortuosity density, central retinal artery/vein equivalents, and local vessel calibre along individual vessel segments in the vessel maps. These vessel segments are defined where crossings and bifurcations are observed.
Supplementary Fig. S7 shows the key differences in global and local vessel calibre. Global vessel calibre is the ratio between vessel pixels and skeletonised vessel pixels (number of yellow pixels divided by blue pixels) (B). This provides a coarse measure of calibre. Local vessel calibre measures the diameter of each vessel segment by traversing it’s length and measuring the diameter of the largest circle which fits entirely within it. A vessel segments’ calibre is then the average value of the list of circle diameters collected as the vessel segment is traversed. Local vessel calibre is thus the overall average of all vessel segment calibres, hence a more granular approach to assess vessel calibre.
Figure S7: SLO binary vessel map (A), binary vessel map with skeleton overlaid in dark blue (B) and binary vessel map with individual vessel segments overlaid in an array of colours (C).
SLOctolyzer’s interface
Supplementary Fig. S8 shows the core steps to run the SLOctolyzer pipeline. SLOctolyzer can be run either from an integrated development environment (IDE), such as Jupyter, Spyder or Visual Studio Code, or directly from the terminal. All that is required is the path to the directory of images to be analysed, and where the results are to be stored in a simple configuration text file. A process log for the user is outputted as the pipeline processes each image in turn, utilising GPU acceleration if available. An optional spreadsheet with additional metadata such as the conversion factor from pixel space to physical space can be provided (in microns-per-pixel), so that measurements in pixel units can be converted into microns.
Figure S8: Summary of the process of running SLOctolyzer on a batch ($n=3$) of images. Directory of input SLO images to be analysed, with an optional spreadsheet which allows manual data entry of pixel length scale, eye type and location (A). Spreadsheet with optional metadata to accompany SLO images during pipeline. Note that missing information is supported (B). Configuration file specifying the file paths to images and where results will be saved (C). Demonstration of running SLOctolyzer via JupyterLab or via the terminal (D).
Supplementary Fig. S9 summarises the output from running SLOctolyzer on a batch ($n=3$) of images. A folder is generated for each SLO image file analysed, and within each folder contains the segmentations masks, feature measurements and a composite image showing the segmentations superimposed onto the SLO. For ease of assessing segmentation quality, composite segmentation images are saved out for every SLO image analysed into a separate folder. Finally, an output spreadsheet collates all feature measurements for every image analysed row-wise, which can be quickly loaded into a statistical programming language for data analysis.
Figure S9: Summary of the output of running SLOctolyzer on a batch ($n=3$) of images. Directory of output folders for each of the three image files analysed (A). Output files for the ‘opticdisc_768’ exemplar SLO image, including the original SLO image, segmentation masks, composite image with segmentations superimposed and a results file with metadata and feature measurements (B). A directory with composite images of SLO and segmentation masks per image file for quick inspection of segmentation quality (C). A composite results file with the metadata and features of each image file processed, stored row-wise to facilitate downstream statistical analysis (D).
Examples of poor performance from reproducibility analysis
Supplementary Fig. S10 shows two examples which had the second largest outliers for tortuosity density and large-vessel calibre (AVR/CRAE/CRVE). In supplementary Fig. S10(A), while there is no qualitative difference in the binary vessel segmentation, tortuosity density is quite large. This is likely due to vessel disconnectedness between the segmentations (purple arrows). Nevertheless, the local vessel calibre and fractal dimension residuals are very low.
Additionally, due to the unregistered nature of the SLO image pairs, minor translations/scales can have an inordinate impact on large-vessel calibre reproducibility given the major arteries and veins appear toward the edge of macula-centred SLO images. In supplementary Fig. S10(B), this particular repeated set of SLO images are off by a horizontal translation while intersecting the disc and thus cuts off one of the major arteries and veins (red arrows). This had a significant consequence to the CRVE and CRAE residuals, causing large error. Nevertheless, local vessel calibre remains stable due to taking into account the whole vasculature of the image.
Figure S10: Some examples of poor agreement in repeated SLO images for tortuosity density (A), and vessel width measurements (B). In Fig. 5, (A) is represented by the blue star, while (B) is represented by the red stars.